Solve (𝐷


2 βˆ’ 3𝐷 + 2)𝑦 = π‘₯


2 + sin π‘₯ where 𝐷 =


𝑑


𝑑π‘₯

1
Expert's answer
2022-01-04T16:48:34-0500

Corresponding homogeneous equation



"(\ud835\udc37^2 \u2212 3\ud835\udc37 + 2)\ud835\udc66 =0"

Characteristic (auxiliary) equation



"r^2-3r+2=0""(r-1)(r-2)=0""r_1=1, r_2=2"

The general solution of the homogeneous differential equation



"y_h=c_1e^x+c_2e^{2x}"

Find the partial solution of the nonhomogeneous differential equation in the form



"y_1(x)=Ax^2+Bx+C+D\cos x+E\sin x""y_1'=2Ax+B-D\sin x+E\cos x""y_1''=2A-D\cos x-E\sin x"

Substitute



"2A-D\cos x-E\sin x-6Ax-3B""+3D\sin x-3E\cos x+2Ax^2+2Bx+2C""+2D\cos x+2E\sin x=x^2+\sin x""x^2: 2A=1""x^1: -6A+2B=0""x^0:2A-3B+2C=0""\cos x:D-3E=0""\sin x: E+3D=1""A=\dfrac{1}{2}, B=\dfrac{3}{2}, C=\dfrac{7}{4}""D=\dfrac{3}{10}, E=\dfrac{1}{10}"

The partial solution of the nonhomogeneous differential equation is



"y_1(x)=\dfrac{1}{2}x^2+\dfrac{3}{2}x+\dfrac{7}{4}+\dfrac{3}{10}\cos x+\dfrac{1}{10}\sin x"

The general solution of the nonhomogeneous differential equation is



"y(x)=c_1e^x+c_2e^{2x}+\dfrac{1}{2}x^2+\dfrac{3}{2}x+\dfrac{7}{4}""+\dfrac{3}{10}\cos x+\dfrac{1}{10}\sin x"